OpenStudy (anonymous):

An arche fires an arrow horizontally with a speed of 89m/s directly at the bull's eye of a target 80 metes away. The arrow, at thr tim it was shot, and the bull;s eye are both 1 meter above the ground. How fa short of the target does the arow strike the ground?

OpenStudy (anonymous):

Okay! So first why don't you try to find out how far away the arrow lands from the archer! Well the time in air can be calculated by using this formula! \[\Delta y = v_{yi} \Delta t + \frac{a \Delta t^{2}}{2}\] But the arrow has no initial vertical component, so that drops out! We know how high it started, and we know it lands at the ground! \[2(y_{f}-y_{i}) = a \Delta t^{2}\] yf is the final position (0) and yi is the intial position 1 \[\sqrt{\frac{2(0 -1)}{a}}= t \] And we know a is -9.8 \[\sqrt{\frac{-2}{-9.8}} = t = .45s\] Okay....then we know that the horizontal distance traveled, is equal to the velocity multiplied by the time in air. \[x = v_{x} \Delta t\] Then just plug in the quantities So it would fall short.... 80 - x